# SpikeLLM: Scaling up Spiking Neural Network to Large Language Models via Saliency-based Spiking

### v1

@@ 5.2 Additive Spiking LLMs
To confirm the additive computation and event-driven sparsity, as in previous binary spiking neural networks, we additionally build $\text{SpikeLLM}_{\texttt{Ter}}$ based on the ternary GIF neuron in Eq.8 as the spiking weight quantizer. We select PB-LLM for comparison since binary weight neural networks (BNNs) can be implemented by accumulating operations. One of the biggest differences between SpikeLLM and PB-LLM is that the spiking neuronal dynamics auto-regressively spike more steps for salient values, instead of the deployment-unfriendly mixed-precision quantization in PB-LLM. In this section, we change the per-channel spiking to unstructured elementwise spiking steps, which is consistent with PB-LLM to improve performance. We further discuss the structured and unstructured spiking issues in Appendix B. In Table 3, we set 3 different spiking steps, where the additional step mask is very small compared to PB-LLM because there are significantly fewer salient weights.
As shown in Table 3, thanks to the spiking dynamics, SpikeLLM gets rid of 8-bit salient outliers in PB-LLM and achieves higher performance. Thus, the fully additive linear layers are achieved. We use the equivalent spiking steps (Table 3) for the simplified operation evaluation and ignore the sparsity, in which the ternary spikes can be more sparse than binary ones via frequency encoding [52]. In Fig.4, SpikeLLM exceeds PB-LLM (BNN) in both effectiveness and efficiency. As the first trial, a large spiking neural network with fully additive linear layers is achieved.
Table 3: Comparisons between SpikeLLM and PB-LLM in LLAMA-2-7B towards additive linear layers. $\text{SpikeLLM}_{\texttt{Ter,x:y:z}}$ indicates using the ternary Integrate-and-Fire neurons in Eq.8 as weight quantizers, where x:y:z is percentages of 1, 2, 4 spiking steps in weights. ACs indicate percentages of computation in linears that implement Multipl-ACcumulates (MACs) as ACcumulates (ACs). We use the Equal Steps as the uniform operation metric, where we view per bit in PB-LLM as an equal step. This is according to both binary- and ternary-weight multiplications are FP16 additions in implementation.
@@
Appendix A Training Details of the OmniQuant Pipeline.
We use a unified training config as shown in all of our experiments. We train LLAMA-1 (7B), LLAMA-2 (7B, 13B, 70B) for both OmniQuant baselines and SpikeLLMs according to this training scheme. Compared with the original OmniQuant, we do not apply loss augmentation methods except for 70B models for training stablity and we set the quantization group number as 1 in all of the experiments. Therefore, this scheme can be viewed as the simplified version without bells and whistles to focus on the influence of the quantization method itself.
In SpikeLLM, we computute the saliency metric for activations layer by layer during the OmniQuant pipeline. Different from activations, we computute the saliency metric for weights directly using the features from the first embedding layer. Because we find this is able to make the computation of inverse Hessian matrix more stable compared with computing layer by layer.
@@
We further investigate the influence of different training samples to confirm the robustness of the proposed methods. In the OmniQuant pipeline, training samples are randomly cropped from the Wikitext2 dataset. To compare the performance of SpikeLLM and the OmniQuant baseline, we use a set of random seeds to sample different training data each time. In Fig. 5, we sample data and train for 40 times, using the same seed for both OmniQuant and SpikeLLM each time. SpikeLLM achieves higher average accuracy on 6 zero-shot datasets and lower Wikitext2 or C4 perplexity at the same time, showing consistently better performance. For the same training data, SpikeLLM always performs better. In other experiments in this paper, we keep the random seed as 2.
Figure 5: Comaprison of different training data for LLAMA-1-7b 4W4A models. In SpikeLLMs, 10% activation channels are set 2 spiking steps.
Appendix B Structured vs. Unstructured Spiking in Weights.
For weight quantization, as shown in Table 5, unstructured spiking steps usually achieve higher performance compared with structured ones. In our per-channel spiking scheme, it can achieve higher performance by setting per-channel spiking steps as elementwise spiking steps. However, unstructured conditions need additional masks and are less friendly to deployment. Therefore, we keep structured settings in low-bit quantization. But for additive LLMs, the performance is more important in the extreme case, and we choose the unstructured settings. Moreover, the PB-LLM baselines are also unstructured, so that, it can also confirm the fair comparison.

### v2

@@ 5.2 Additive Spiking LLMs
To direct maintain the addition nature and event-driven sparsity of SNNs without expanding GIF neurons, we additionally build $\text{SpikeLLM}_{\texttt{Ter}}$ based on the ternary GIF neuron in Eq.6 as the spiking weight quantizer. After weight quantization as Eq.6, linear layers can be implemented by ACcumulation (AC). The detailed training and evaluation settings are reported in Appendix A.3.
Comparison with binary LLM. We select PB-LLM for comparison since binary weight neural networks (BNNs) can be implemented by ACs. As shown in Table 6, SpikeLLM achieves full AC operations in linear layers by saliency-based spiking neuronal dynamics, instead of the deployment-unfriendly mixed-precision quantization in PB-LLM.
Efficiency of additive SpikeLLM. In Table 6, we evaluate the AC operations in a linear layer by equal-steps. For PB-LLM, we view the average bit-width as the equal steps; for SpikeLLM, the equal steps are calculated by the average spiking steps of salient and other values. As shown in Table 6, SpikeLLM is able to exceed PB-LLM with similar equal-steps. Further, in Fig 5, we evaluate the Pareto front about equal-steps and performance, which shows SpikeLLM exceeds PB-LLM (BNN) in both effectiveness and efficiency.
Table 6: Comparisons between SpikeLLM and PB-LLM in LLAMA-2-7B towards additive linear layers. $\text{SpikeLLM}_{\texttt{Ter,x:y:z}}$ indicates using the ternary GIF neurons in Eq.6 as weight quantizers, where x:y:z is percentages of 1, 2, 4 spiking steps in weights (details in Appendix A.3).
@@
6 Conclusion
We propose the first spiking large language models with 7$\sim$70 billion parameters, promoting SNNs to the era of LLMs. This is achieved by significant improvement of spike encoding efficiency, where the GIF neurons compress the code length from T to $\frac{T}{L}log_{2}L$ bits; the OBSpiking further compresses to approximate $log_{2}T$ bits. Unlike previous ANN-SNN conversions that rely on quantization, this work exceeds quantization with the hybrid encoding of both SNNs and quantized-ANNs. Scaling up spiking neuronal dynamics in LLMs and developing training efficient spiking mechanisms for tens of billion parameters have become increasingly valuable issues.
Limitations A primary limitation of this paper is the performance gap between Spiking LLMs and ANN LLMs. This work applies ANN-SNN calibration with only 128 calibration data and finishes training a 7B model in 3.5 GPU hours. Continuous pretraining would boost the performance.
Acknowledgments
This work is supported by National Science and Technology Major Project (2023ZD0121504). This work is partially supported by CAS Project for Young Scientists in Basic Research (YSBR-116), National Natural Science Foundation of China (62325603, 62236009), Beijing Science and Technology Plan (Z241100004224011).
Appendix A Appendix
A.1 Low-Bit Quantization
This indicates the complexity and energy consumption of MAC operations are proportional to $M\times N$. Therefore, the number of operations in a quantized model can be measured using the arithmetic computation effort (ACE) metric (Zhang et al., 2022b), which is defined as $M\times N$ for a MAC operation between the M-bit weight and N-bit activation. Recent LLMs contain more than 10B $\sim$ 100B parameters, making it an essential requirement to push not only weights but also activations to lower bit-width.
A.2 Training Details of the OmniQuant Pipeline.
A.2.1 Training Details.
We use a unified training config as shown in all of our experiments. We train LLAMA-1 (7B), and LLAMA-2 (7B, 13B, 70B) for both OmniQuant baselines and SpikeLLMs according to this training scheme. Compared with the original OmniQuant, we do not apply loss augmentation methods except for 70B models for training stability and we set the quantization group number as 1 in all of the experiments. Therefore, this scheme can be viewed as a simplified version without bells and whistles to focus on the influence of the quantization method itself.
In SpikeLLM, we compute the saliency metric for activations layer by layer during the OmniQuant pipeline. Different from activations, we compute the saliency metric for weights direct using the features from the first embedding layer. Because we find this can make the computation of the inverse Hessian matrix more stable compared with computing layer by layer.
@@ A.1 Low-Bit Quantization
A.2.2 Training Data.
Following OmniQuant (Shao et al., 2023), we randomly select 128 calibration data from WikiText2, which are 2048-token chunks. We further investigate the influence of different training samples to confirm the robustness of the proposed methods. In the OmniQuant pipeline, training samples are randomly cropped from the Wikitext2 dataset. To compare the performance of SpikeLLM and the OmniQuant baseline, we use a set of random seeds to sample different training data each time. In Fig. 6, we sample data and train 16 times, using the same seed for both OmniQuant and SpikeLLM each time. SpikeLLM achieves higher average accuracy on 6 zero-shot datasets and lower Wikitext2 or C4 perplexity at the same time, showing consistently better performance. For the same training data, SpikeLLM always performs better. In other experiments, we keep the random seed as 2.
Figure 6: Comparison of different training data for LLAMA-1-7b W4A4 models. In SpikeLLMs, 10% activation channels are set to 2 spiking steps. Loss augmentations are applied as OmniQuant (Shao et al., 2023). The average accuracy (not norm) is evaluated.
A.3 Training Details of the Ternary SpikeLLM.
A.3.1 Model Definition.
In additive SpikeLLM, we apply the ternary spike level {-1,0,+1} to encode weights according to Eq. 6. After weight ternarization, the matrix multiplication in the linear layer can be implemented by full ACcumulation (AC). As shown in Table 8, we have 3 spiking-step settings. For example, for the $\text{SpikeLLM}_{\texttt{Ter,70:25:5}}$ model, we allocate the most salient 5% values in weights with 4 spiking steps, the following 25% with 2 spiking steps, and the rest 70% with 1 spiking steps. The unstructured spiking steps are applied, which is different from the experiment in the Main Result Section. As the following section, this is because of more accurate quantization in ultra-low bit quantization and fair comparison with PB-LLM (Shang et al., 2023).
Table 8: Model settings in additive SpikeLLM. 1,2,4-Steps indicate the percentages of 1, 2, and 4 spiking-step encoding in weights respectively.
Model 1-Step 2-Step 4-Step Avg. Step Spike-Level
$\text{SpikeLLM}_{\texttt{Ter,70:25:5}}$ 70% 25% 5% 1.4 {-1,0,+1}
$\text{SpikeLLM}_{\texttt{Ter,80:15:5}}$ 80% 15% 5% 1.3 {-1,0,+1}
$\text{SpikeLLM}_{\texttt{Ter,85:10:5}}$ 85% 10% 5% 1.25 {-1,0,+1}
$\text{SpikeLLM}_{\texttt{Ter,90:5:5}}$ 90% 5% 5% 1.2 {-1,0,+1}
A.3.2 Structured vs. Unstructured Spiking in Weights.
For weight quantization, as shown in Table 9, unstructured spiking steps usually achieve higher performance compared with structured ones. Our per-channel spiking scheme can achieve higher performance by setting per-channel spiking steps as elementwise spiking steps. However, unstructured conditions need additional masks and are less friendly to deployment. Therefore, we keep structured settings in low-bit quantization. But for additive LLMs, the performance is more important in the extreme case, and we choose the unstructured settings. Moreover, the PB-LLM baselines are also unstructured, so that, it can also confirm the fair comparison.
@@
A.3.3 Training Details.
The same as PB-LLM (Shang et al., 2023), we follow the GPT-Q (Frantar et al., 2022) pipeline for the post-training weight quantization. In quantization, we keep the same block size of 128. We apply the same 128 calibration data as PB-LLMs. Each data is randomly selected from WikiText2 and tokenized to formulate a 2048-token chunk.
A.4 Spiking Neuronal Dynamics
A.4.1 Step Merging in GIF Neurons
In Fig. 7, we illustrate the function Eq. (5). After merging multiple steps (L) of the IF neuron, the multi-step spike become a floor function, which can activate multi-level firing. The membrane potential of the GIF neuron will fire according to Eq. (5). Finally, the output of GIF neuron is according to Eq. (3) : $\mathbf{x}^{(\ell-1)}(t)=\mathbf{s}^{(\ell)}(t)V_{th}+\min{(\mathbf{x}^{(\ell-1)})}$.
Figure 7: The floor firing function of the GIF neuron.
Figure 8: Neuronal dynamics of GIF neurons. We compare the neuron input, membrane potential, and output spike in three conditions: (a) the traditional IF neuron. (b) Multi-step merged GIF neuron. (c) Binary inference of GIF neuron.
Figure 9: Training dynamics of GIF neurons. We compare the traditional backpropagation through time (BPTT), quantization-aware training in GIF neurons.
A.4.2 Comparison of Inference
(i) Neuronal Dynamics of IF Neurons (Fig. 8 (a)): Given a real-valued input sequence, the membrane potential accumulates over time. When it exceeds the firing threshold, the IF neuron emits a spike, and the membrane potential is reset by subtracting the threshold value.
(ii) Neuronal Dynamics of Multi-Step GIF Neurons (Fig. 8 (b)): Similar to IF neurons, GIF neurons accumulate membrane potential over time steps. However, unlike IF neurons, GIF neurons first integrate inputs over a fixed number of steps L. Based on Eq. (5), GIF neurons can fire k discrete levels of spikes, where $k\in[0,L]$.
(iii) Neuronal Dynamics of Binary GIF Neurons (Fig. 8 (c)): The only difference between multi-bit and binary inference of GIF neurons lies in the firing threshold. GIF neurons in multi-bit inference have multiple thresholds corresponding to different firing levels, while binary inference uses a single threshold, similar to IF neurons. As a result, binary inference emits one bit of spike per time step, as a result, multi-level spikes are effectively decoded as binary. Next, we demonstrate the equivalence of binary inferenced GIF neurons and IF neurons.
Equivalence Conditions for Spiking Neurons:
(i) Given the same real-valued input sequence with firing rate encoding, the neurons are considered equivalent if the spike firing rates are identical.
(ii) For finite-step spikes, the membrane potential at the end of a given period must also be the same to ensure consistent firing rates in subsequent steps.
Based on these conditions, since the binary GIF and IF neurons share the same inputs and firing threshold over a short time window, their firing rates and membrane potential values are identical, confirming their equivalence in binary inference.
A.4.3 Comparison of Training
As shown in Fig. 9, we compare the training dynamics of conventional spiking neural networks (SNNs) and GIF neurons. The discussions are as follows.
(i) Fig. 9 (a): Traditional backpropagation through time (BPTT) for IF (or LIF) neurons suffers from significant quantization errors due to binary quantization at each time step. These errors accumulate over time, making SNNs difficult to train.
(ii) Fig. 9 (b): GIF neurons use quantization-aware training. By aggregating L time steps, GIF neurons perform multi-bit quantization, which eliminates inter-step error accumulation. The multi-bit quantization improves gradient estimation (STE), leading to more accurate training.
A.5 More Results
A.5.1 Perplexity in OmniQuant Settings
Table 10: Comparisons between SpikeLLM and OmniQuant in the Wikitext2 and C4 PPL metrics. The LLAMA-2 family is evaluated. We do not evaluate the W4A4 and W2A8 settings for LLAMA-2-70B because the grouped-query attention (GQA) makes training unstable in the OmniQuant pipeline.
LLAMA-2-7B LLAMA-2-13B LLAMA-2-70B
Method Saliency #Bits Wikitext2 C4 Wikitext2 C4 Wikitext2 C4
OmniQuant W4A4 15.25 19.35 12.40 15.87
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W4A4 11.93 15.34 9.64 12.48
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W4A4 11.85 15.18 12.17 15.71
OmniQuant W2A8 287.64 445.21 53.87 72.33
$\text{SpikeLLM}_{\texttt{T=4}}$ 0.05 W2A8 29.98 46.61 12.80 17.06
OmniQuant W2A16 38.05 98.74 17.14 27.12 10.04 19.31
$\text{SpikeLLM}_{\texttt{T=2}}$ 0.10 W2A16 14.16 19.73 9.45 13.86 6.12 8.82
A.5.2 Comparison with QLoRA Methods
This line of work focuses on adding a low-rank branch to finetune the low-bit quantized model. State-of-the-art methods including IR-QLoRA (Qin et al., 2024), QA-LoRA (Xu et al., 2023), QLoRA (Dettmers et al., 2024). Before training, we calibrate saliency and acquire a 2-bit weight spiking model with the EfficientQAT (Chen et al., 2024) setting. We follow the same settings of QLoRA and IR-LoRA in Alpaca finetuning.
Table 11: Performance comparison in QLoRA Settings.
Method Data #Bit Hums. STEM Social Other Avg.
LLAMA-7B - 16 33.3 29.8 37.8 38.0 34.6
NormalFloat - 2 24.2 28.9 31.1 25.0 26.9
QLoRA (w/ GPTQ) Alpaca 2 23.4 26.2 26.4 28.4 25.8
QLoRA Alpaca 2 24.0 27.0 27.5 26.7 26.2
QA-LoRA Alpaca 2 27.3 26.1 26.1 30.3 27.5
IR-QLoRA Alpaca 2 26.0 27.8 30.2 28.3 27.8
[…]
A.5.3 Comparison in Efficient QAT Settings
For low-bit weight-only quantizations, we train and evaluate SpikeLLM following EfficientQAT (Chen et al., 2024) (including training data and evaluation metrics, lm-eval-v0.4.2). Both the EfficientQAT and SpikeLLM are trained in 2048 length in two stages with the same condition. We compare the 2-bit performance as follows.
Table 12: Performance comparison in Efficient QAT pipeline.
Method #Bit WinoGrande HellaSwag ArcC ArcE PiQA Avg.
LLAMA-2-7B - 69.22 57.18 43.17 76.09 78.24 64.78
GPTQ 2 55.17 32.59 21.25 40.45 58.32 41.56
OmniQ 2 55.88 40.28 23.46 50.13 65.13 46.98
AutoRound 2 61.01 40.28 32.25 65.99 72.96 54.50
AQLM2×8 2 65.27 49.96 32.85 66.92 73.07 57.61
EfficientQAT 2 65.19 49.53 35.15 69.74 73.67 58.65
[…]